You are considering the purchase of a simple insurance policy on your home. The annual premium...

  1. You are considering the purchase of a simple insurance policy on your home. The annual premium is $1,500. For this simplified example, let’s assume that the probability of no loss occurring during the policy year is 95%. In that case, your net financial benefit is the full premium amount, so the net impact is -$1,500. The probability of a small loss ($1,000) is 3%, in which case your net financial impact would be -$500 (the premium you paid plus the loss the policy covers) (-$1,500 + $1,000). The probability of a moderate loss ($10,000) is 1.8%. The probability of a total loss ($400,000) is 0.2%.

What is the expected value of the net financial impact for you? (show calculations)

Would you buy a policy like this? Why?

How is this different from the raffle?

For the insurance company, the financial impact is the premium they collect from you and many other policyholders minus the losses minus their expenses (around 1/3 of the premium). Why might their perspective be different from yours?.
Answer #1

Answer:-

Given That:-

the purchase of a simple insurance policy on your home. The annual premium is $1,500. For this simplified example, let’s assume that the probability of no loss occurring during the policy year is 95%. In that case, your net financial benefit is the full premium amount, so the net impact is -$1,500. The probability of a small loss ($1,000) is 3%, in which case your net financial impact would be -$500 (the premium you paid plus the loss the policy covers) (-$1,500 + $1,000). The probability of a moderate loss ($10,000) is 1.8%. The probability of a total loss ($400,000) is 0.2%.

\rightarrow The given value are table from and calculation is shown below:

\rightarrow

loss 0 $1.000 $10,000 $400,000
probability 0.95 0.03 0.018 0.002

\rightarrow and. the loss to the converted premium

premium $1,500 $1,500 $1,500 $1,500
net impact -1,500 -5,00 +8500 +398,500

now we are find out is

\rightarrowThe expected value of the probability of.a total loss.?

Here we know the formula \sum (net impact \times probability)

So, now substitute the values

=(0.8\times 1500)+(0.03\times -500)+(0.018\times 8500)+(0.002\times 398,50)

=(-1200)+(-15)+(153)+(797)

=-265

expected value of the "p" of a total loss pl -265

*\rightarrow The buy a policy like their ?why?

\rightarrow This is negative financial input , that why there is not buy a policy

*\rightarrow The might their perspective be different from your?

\rightarrowThere pl the also have the negative net financial in pact , is her wont's will be different

\rightarrowIt is the need the minimize the amount to they have to pay to policy holders to increase profit.

\rightarrowSo, the will pass to buy this  policy